### v2

# Chaos and spatial prethermalization in driven-dissipative bosonic chains

Abstract
Thermalization in quantum many-body systems, the process by which they naturally evolve toward thermal equilibrium, typically unfolds over timescales set by the underlying relaxation mechanisms. Yet, the spatial aspect of thermalization in these systems is less understood. We investigate this phenomenon within the nonequilibrium steady state (NESS) of a Bose-Hubbard chain subject at its boundaries to coherent driving and dissipation, a setup inspired by current designs in circuit quantum electrodynamics. We uncover a two-stage thermalization process along the spatial dimension. Close to the coherent drive, the U(1) symmetry of the phase of the photonic field is restored over a short length scale, while its amplitude relaxes over a much larger scale. This opens up an extensive region of the chain where the photon density remains high, and the chaotic dynamics give rise to a hydrodynamic regime, characterized by local equilibria with a large and slowly-varying effective chemical potential. Dynamical fingerprints of chaos in this NESS are probed using semiclassical out-of-time-order correlators (OTOCs) within the truncated Wigner approximation (TWA). We explore the conditions underlying this protracted thermalization in space and argue that similar prethermal chaotic phases are likely to occur in a broad range of extended driven-dissipative systems.
I Introduction
Thermalization underpins statistical mechanics: generic isolated interacting systems evolve towards thermal equilibrium regardless of their initial conditions. Classically, this is addressed by Boltzmann’s H-theorem which relies on the assumption of molecular chaos [1, 2]. Quantum mechanically, signatures of chaos manifest themselves in the structure of the eigenvalue spectrum and the Eigenstate Thermalization Hypothesis provides a theoretical framework to explain how closed Hamiltonian systems can achieve thermalization under unitary dynamics [3, 4, 5, 6].
Typically, thermalization occurs in two stages. First, through a rapid relaxation of non-conserved quantities to local equilibrium values. Then, the hydrodynamic modes—long-wavelength excitations associated with conserved quantities in the bulk of the system such as energy or particle density—relax by a slower, often diffusive, process. The Fermi-Pasta-Ulam-Tsingou paradox—the apparent lack of thermalization in a nonlinear chain of classical oscillators—posed an unexpected challenge to this understanding  [7]. It is now widely accepted that this system’s proximity to an integrable point is responsible for prolonged transient dynamics [8]. In quantum mechanics, long-lived transients are similarly observed in near-integrable systems. When an integrable system, characterized by many conserved quantities, undergoes a quench in which an integrability-breaking pertubation is suddenly turned on, it initially relaxes to a quasi-equilibrium state described by a Generalized Gibbs Ensemble [9, 10, 11, 12, 13]. The subsequent relaxation to a true thermal state can be exceedingly slow, especially when the integrability-breaking interaction is weak [14]. Such slow thermalization dynamics are also seen in periodically-driven systems, where high-frequency driving lead to the emergence of long-lived “prethermal” states [15, 16, 17, 18, 19]. These states retain memory of the initial conditions through approximately conserved quantities, and theoretical proposals and experiments have been put forward to leverage these time-dependent protocols for quantum state engineering [20, 21, 22, 23, 24].
Figure 1: Schematic of the boundary-driven dissipative Bose-Hubbard chain and key results. (a) Tight-binding array of $L$ nonlinear resonators described by the Hamiltonian in Eq. (2) subject to a drive of amplitude $F$ coherently injecting photons at the leftmost site and to single-photon losses at both ends of the chain. The interplay of interaction, drive, and dissipation leads to a nonequilibrium steady state (NESS). (b) Lyapunov growth between neighboring Wigner trajectories are used to identify chaotic dynamics. (c) In the chaotic regime, the chain hosts three distinct domains illustrated by their steady-state local Wigner functions $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$: a nonsymmetric nonthermal domain near the left boundary, an extensive prethermal phase where the U(1)-symmetry of the phase is restored and which hosts a high density of photons only saturated by the Kerr non-linearity, and a U(1)-symmetric thermal domain near the right boundary which is characterized by thermal fluctuations over the vacuum state.
With recent advances in quantum device fabrication, these theoretical considerations have now gained significant experimental relevance for quantum technologies and can now be addressed by specifically engineering many-body systems. In particular, in circuit quantum electrodynamics, nonlinear oscillators are routinely constructed by coupling microwave resonators to nonlinear elements such as Josephson junctions. These building blocks can be assembled into long chains of nonlinear oscillators with up to nearly 100 sites, as demonstrated in studies such as Refs. [25, 26, 27], and they can be effectively described by Bose-Hubbard models. Similar chains have been realized in other quantum architectures, including trapped ions [28], semiconductor micropillars [29], and ultracold gases in optical lattices [30].
These devices are inherently prone to intrinsic losses as well as extrinsic dissipation channels introduced by measurement apparatuses. Furthermore, they are operated away from thermal equilibrium by processes such as gate sequences in quantum computers, pulsed signals in quantum optimal control, or continuous wave drives in quantum optical setups. This open, nonequilibrium nature poses significant conceptual challenges: established concepts such as chaos, thermalization, and the potential for persistent prethermal states must be reconsidered in this context of driven-dissipative dynamics. Pioneering works with classical systems, such as the boundary-driven dissipative Klein-Gordon chain, have demonstrated that the interplay between drive, dissipation, and interactions can create a complex dynamical landscape, including unconventional transport phenomena [31, 32, 33], which are simply absent in their isolated counterparts [34].
In this work, we investigate the route to thermalization in quantum chains of nonlinear bosonic modes. Photons are coherently injected at one end of the chain, and the coupling to the input feedline and output channels brings incoherent photon losses at both ends of the chain. We leverage the fact that these systems operate in regimes well described by semiclassical approaches to promote well-established concepts and tools from classical chaos theory to analyze the ergodic properties of the nonequilibrium steady state (NESS) of the open quantum chain. Specifically, we use the truncated Wigner approximation (TWA) and out-of-time-order correlators (OTOCs) to follow the transition from regular to chaotic regimes. In the latter, we examine the thermalization along the chain, spatially resolving the transition from the strongly nonequilibrium state near the coherent drive to the expected thermal state at the opposite end of the chain. Between these extremes, we identify an extended prethermal phase where the U(1) symmetry, initially broken by the boundary drive, is restored and the chain locally equilibrates to high-density states characterized by a large and slowly-varying effective chemical potential. We attribute this phenomenon to a significant mismatch between the short relaxation scale of the phase degree of freedom and the longer hydrodynamic relaxation of the amplitude sector of the photonic field. Our findings are directly relevant to current experimental platforms, and we propose diagnostics based on routinely measured quantities, which can be determined through quantum-state tomography via, e.g., heterodyne detection.
II Results
II.1 Boundary-driven dissipative Bose-Hubbard chain
We consider a one-dimensional chain of $L$ single-mode photonic resonators with nearest-neighbor coupling, modeled by the Bose-Hubbard model. The leftmost resonator of the chain is driven by a continuous wave drive, and the resonators at both ends of the chain experience single-photon losses, as depicted in Fig. 1. The intrinsic losses of the resonators within the bulk of the chain are assumed to be negligible. The dynamics are modeled by a Lindblad master equation for the system’s density matrix $\hat{\rho}(t)$ reading (we set $\hbar=1$)
$\frac{\partial\hat{\rho}}{\partial t}=-{\rm i}[\hat{H},\hat{\rho}]+\mathcal{D}[\hat{L}_{1}]\hat{\rho}+\mathcal{D}[\hat{L}_{L}]\hat{\rho}\,.$ (1)
The system Hamiltonian is expressed in the frame rotating at the drive frequency $\omega_{\rm d}$ as
$\hat{H}=$ $\sum_{\ell=1}^{L}\left(-\Delta\,\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}+\frac{1}{2}\,U\,\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\hat{a}_{\ell}\right)$ (2)
$-J\sum_{\ell=1}^{L-1}\big{(}\hat{a}_{\ell+1}^{\dagger}\hat{a}_{\ell}+\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell+1}\big{)}+F(\hat{a}_{1}^{\dagger}+\hat{a}_{1})\,.$
Here, $\hat{a}_{\ell}^{\dagger}$ ($\hat{a}_{\ell}$) are the bosonic creation (annihilation) operators for the photons in the $\ell$-th resonator mode of frequency $\omega_{0}$. $\Delta:=\omega_{\rm d}-\omega_{0}$ is the pump-to-resonator detuning, $J$ is the hopping amplitude between neighboring resonators, and $U$ is the strength of the onsite Kerr nonlinearity. Our model is motivated by state-of-the-art experimental cavity or circuit QED setups [25, 26, 27] with a weak but finite interaction, $|U|\ll|\Delta|,|J|$. While $U=0$ makes the model trivially integrable, a finite $U$ ensures the nonintegrability of the Bose-Hubbard Hamiltonian on sizeable chains [35]. $F$ is the amplitude of the driving field, which explicitly breaks the U(1)-symmetry of the Bose-Hubbard chain. We take $U>0$, $\Delta>0$, $J>0$, and $F\geq 0$ ($U<0$, $\Delta<0$, $J<0$, and $F\leq 0$ yields identical results).
The Lindblad dissipators at sites $\ell=1$ and $\ell=L$ are defined as
$\mathcal{D}[\hat{L}_{\ell}]\hat{\rho}:=\hat{L}_{\ell}\hat{\rho}\hat{L}_{\ell}^{\dagger}-\frac{1}{2}\left\{\hat{L}_{\ell}^{\dagger}\hat{L}_{\ell},\hat{\rho}\right\},$ (3)
where $\hat{L}_{\ell}=\sqrt{\gamma}\hat{a}_{\ell}$ are the local jump operators modeling incoherent one-photon losses to the environment at a rate $\gamma>0$. When nondriven and isolated, $F=\gamma=0$, the Bose-Hubbard chain is far from the Mott insulating regime, and it is naturally expected to thermalize (see, e.g.[36, 37]). In the presence of drive and dissipation, it is expected to reach a unique NESS, $\hat{\rho}^{\rm ss}:=\lim\limits_{t\to\infty}\hat{\rho}(t)$, carrying a uniform DC current of photons from the left drive to the right drain. The initial conditions to Eq. (1) are inconsequential to our results and, unless stated otherwise, we simply use the vacuum state $\hat{\rho}(0)=\bigotimes_{\ell=1}^{L}|0\rangle_{\ell}\langle 0|_{\ell}$ where $|0\rangle_{\ell}$ is the Fock state with no excitation in the $\ell$-th resonator. We take $\gamma$ as the unit of energy and set $\Delta=2.5$, $J=2$, and $U=0.1$ for the remainder of this study.
To access the NESS of Eq. (1), we use the truncated Wigner approximation (TWA), an approach based on a semiclassical treatment of the bosonic fields that accounts for leading-order quantum fluctuations [38, 39]. In the TWA, Eq. (1) is mapped to a set of $L$ coupled stochastic differential equations for the complex field amplitudes $\alpha_{\ell}$, reading
${\rm i}\frac{\partial\alpha_{1}}{\partial t}$ $=-f(\alpha_{1})-J\alpha_{2}+F-\frac{{\rm i}\gamma}{2}\alpha_{1}+\sqrt{\frac{\gamma}{2}}\xi_{1}(t)\,,$
${\rm i}\frac{\partial\alpha_{\ell}}{\partial t}$ $=-f(\alpha_{\ell})-J(\alpha_{\ell-1}+\alpha_{\ell+1})\,,\ l=2,...,L-1$
${\rm i}\frac{\partial\alpha_{L}}{\partial t}$ $=-f(\alpha_{L})-J\alpha_{L-1}-\frac{{\rm i}\gamma}{2}\alpha_{L}+\sqrt{\frac{\gamma}{2}}\xi_{L}(t)\,,$ (4)
[…]
Although the TWA is exact only for quadratic models ($U=0$), it has been successfully applied to describe dissipative phase transitions, disordered systems, time crystals, and quantum chaos in a variety of weakly nonlinear driven-dissipative systems [40, 41, 42, 43]. We justify the use of the TWA in our model by the weakness of the Kerr nonlinearity—a condition easily achievable in current circuit QED architectures [44]; under this condition, the TWA faithfully describe the NESS of single or multiple nonlinear driven resonators [40] and boundary-driven dimers [43]. We further validate this approach by benchmarking it against exact results for small system sizes and comparing it with other approximation schemes (see the Supplementary Information). The choice of the TWA is motivated by the considerable reduction of the computational complexity it offers: the exponential growth of the Hilbert space in Eq. (1) is cut down to a linear growth in the number of the stochastic equations (II.1), thus enabling direct numerical integration of long chains of resonators up to times past the transient dynamics and deep in the steady-state regime. Furthermore, Wigner trajectories extend the classical notion of phase-space trajectories to the quantum regime, thereby offering a practical means to investigate the exponential sensitivity to initial conditions (or the lack thereof), which is a constitutive hallmark of chaotic dynamics [45]. Notably, while the density matrix and the local Wigner functions remain constant in the NESS, individual Wigner trajectories keep fluctuating and can thus be used to analyze the integrable versus chaotic character of the dynamics in the NESS [46, 47].
Figure 2: Steady-state quantum chaos via phase OTOC. Dynamics of the steady-state phase OTOC $D_{1,\ell}^{\rm ss}(\tau)$, defined below Eq. (5), computed at sites $\ell=1,10,20,...,80$ of an $L=80$ chain and for two representative drive strengths $F$. (a) $F=7.5$: Exponential growth at a Lyapunov rate $\lambda\approx 2.8$ and ballistic spreading at a butterfly velocity $v=2J$, signaling chaotic dynamics. (b) $F=12$: Sub-exponential growth, signaling non-chaotic or regular dynamics. Dashed lines indicate the chaotic bound on the saturation value of the phase OTOC. The results are computed upon averaging over $5\times 10^{3}$ independent Wigner trajectories. Throughout the manuscript, the dissipation rate $\gamma$ is taken as the unit of energy unit and the other parameters are set to $\Delta=2.5$, $J=2$, and $U=0.1$.
II.2 Regular and chaotic regimes in the NESS
Dynamical signatures of quantum chaos are commonly probed by means of out-of-time-order correlation functions such as average square commutators of the form $\textrm{Tr}\left([\hat{W}_{\ell}(\tau),\hat{V}_{k}]^{\dagger}[\hat{W}_{\ell}(\tau),\hat{V}_{k}]\hat{\rho}\right)$ where $\hat{V}_{k}$ and $\hat{W}_{\ell}$ are operators with local support around sites $k$ and $\ell$, respectively [48, 49, 50]. We focus on the OTOC between the number and the phase degrees of freedom of the resonators, i.e., $\hat{n}_{k}$ and $\hat{\phi}_{\ell}$ evaluated at time $t$ and later time $t+\tau$, respectively. In the Methods, we show that the semiclassical formulation of this phase OTOC can be computed as [51, 52, 53, 54, 55, 56, 57]
$D_{k,\ell}(t,\tau):=1-\left\langle\cos\left[\phi_{\ell}^{(a)}(t+\tau)-\phi_{\ell}^{(b)}(t+\tau)\right]\right\rangle.$ (5)
Here, $\phi_{\ell}(t):=\textrm{arg}\,\alpha_{\ell}(t)$ is the phase of the complex field in the $\ell$-th resonator along an individual Wigner trajectory. The superscripts $(a)$ and $(b)$ refer to two replicas of the system which are identical until time $t$ at which an infinitesimal local perturbation is applied at site $k$ in replica $b$: $\alpha_{k}^{(b)}(t)\to\alpha_{k}^{(b)}(t)+\varepsilon$. The subsequent evolution is computed using the same quantum noise realization for the two replicas. delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the average over realizations of the noise and the limit $\varepsilon\to 0$ is implied.
Figure 3: Nonequilibrium steady-state phase diagram. Steady-state properties of the last resonator in the chain, $\ell=L$, as a function of the chain length $L$ and drive strength $F$. (a) Saturation value of the steady-state phase OTOC, $D^{\rm ss}_{1,L}(\tau\to\infty)$, defined below Eq. (5). (b) Photon number $n^{\rm ss}_{L}$. (c) Photon-number fluctuation $\delta n_{L}$ defined in Eq. (6). (d-f) Normalized local Wigner functions, $\tilde{W}^{\rm ss}_{L}(\alpha,\alpha^{*}):=W_{L}^{\rm ss}(\alpha,\alpha^{*})/\textrm{max}[W^{\rm ss}_{L}(\alpha,\alpha^{*})]$, for chain lengths $L=10$ (red colorscale) and $L=80$ (blue half-width contours) and for drive strengths (d) $F=2$, (e) $F=7.5$, and (f) $F=12$, corresponding to the markers in panel (a). The three distinct regimes labelled by (I), (II) and (III) are discussed in the text. Results are computed upon averaging over $10^{3}$ independent Wigner trajectories. Statistics are further improved by averaging over a time window $\Delta\tau$ after the NESS is reached, with $\Delta\tau=25$ for $D^{\rm ss}_{1,L}(\tau\to\infty)$, and $\Delta\tau=500$ for $n^{\rm ss}_{L}$ and $\delta n^{\rm ss}_{L}$. The other parameters are set as in Fig. 2.
We compute the steady-state phase OTOC, defined as $D^{\rm ss}_{k,\ell}(\tau):=\lim\limits_{t\to\infty}D_{k,\ell}(t,\tau)$, to map the nonequilibrium steady-state phase diagram of the boundary-driven dissipative Bose-Hubbard chain. This quantity successfully captures basics features of quantum information spreading in extended many-body systems [58] and we show in the Methods that $D^{\rm ss}_{1,\ell}(\tau)$ clearly exhibits a causal light-cone structure with a ballistic spreading of information characterized by a butterfly velocity $v=2J$ [59]. On general grounds, we have $D_{k,\ell}^{\rm ss}(\tau=0)=0$, $D_{k,\ell}^{\rm ss}(\tau>0)$ is expected to increase with $\tau$ whenever trajectories in the two replicas start deviating, and, eventually, to saturate to a finite value $D_{k,\ell}^{\rm ss}(\tau\to\infty)$. Both the growth and the saturation regimes of $D_{k,\ell}^{\rm ss}(\tau)$ shed light on the chaotic versus regular nature of the dynamics in the NESS. In chaotic dynamics, the typical distance between two trajectories with nearly identical initial conditions is expected to grow exponentially – a hallmark of Lyapunov instability, which is often taken a defining feature of classical chaos. Quantum mechanically, this exponential growth is expected to be captured by OTOCs in systems that have a well-defined semiclassical limit or large-$N$ limit. The saturation value $D_{k,\ell}^{\rm ss}(\tau\to\infty)$ is related to the square of the available local phase space volume. $D_{k,\ell}^{\rm ss}(\tau\to\infty)\approx 0$ corresponds to situations where the phases in replica $a$ and $b$ remain strongly correlated, suggestive of regular dynamics. In contrast, in chaotic regimes where the trajectories decorrelate rapidly, $\phi^{(a)}$ and $\phi^{(b)}$ can be seen as uniformly distributed random phases, yielding a saturation value at its chaotic bound, $D^{\rm ss}_{k,\ell}(\tau\to\infty)\approx 1$.
In Fig. 2, we illustrate the growth of the steady-state phase OTOC for two representative values of the drive strength $F$. In panel (a), $D^{\rm ss}_{1,\ell}(\tau)$ shows a rapid exponential growth of the form $D^{\rm ss}_{1,\ell}(\tau)\sim\exp[\lambda(t-l/v)]$ where $\lambda$ is a Lyapunov rate, followed at scrambling times on the order of $1/2J$ by a saturation regime where $D^{\rm ss}_{1,\ell}(\tau\to\infty)\approx 1$, i.e., maximal decorrelation. Altogether, the semiclassical phase OTOC $D_{1,\,\ell}^{\rm ss}(\tau)$ captures both the Lyapunov growth and the saturation regimes expected of quantum chaotic dynamics. In panel (b), $D^{\rm ss}_{1,\ell}(\tau)$ instead displays early-time oscillations and the overall growth is slower than exponential. Eventually, the late-time saturation value $D^{\rm ss}_{1,\ell}(\tau)$ is significantly less than $1$. This is indicative of regular dynamics.
Beyond the two representative values of $F$ used in Fig. 2, we verified across the examined parameter range that $D^{\rm ss}_{1,\ell}(\tau\to\infty)=1$ consistently corresponds to exponential growth, while $D^{\rm ss}_{1,\ell}(\tau\to\infty)\ll 1$ corresponds to sub-exponential growth. Therefore, we use the saturation value of the steady-state phase OTOC, $D^{\rm ss}_{1,\ell}(\tau\to\infty)$, as a proxy to map the chaotic versus regular character of the nonequilibrium steady-state dynamics as the drive strength $F$ and the chain length $L$ are varied. Additionally, we monitor two standard quantities in quantum optics, namely the steady-state photon number $n_{\ell}^{\rm ss}:=\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle_{\rm ss}$ and its fluctuations $\Delta n_{\ell}^{\rm ss}:=\langle(\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell})^{2}\rangle_{\textrm{ss}}-\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle_{\textrm{ss}}^{2}$ by means of the quantity
$\delta n^{\rm ss}_{\ell}:=\Delta n^{\rm ss}_{\ell}-n^{\rm ss}_{\ell}$ (6)
which quantifies the relative distance to Poissonian statistics  [^1] Compared to the conventional indicator $g^{(2)}$, $\delta n^{\rm ss}_{\ell}$ defined in Eq. (6) is less prone to numerical errors caused by a finite sampling of Wigner trajectories.. $\langle...\rangle_{\rm ss}$ indicates the average once the NESS is reached. $\delta n^{\rm ss}_{\ell}=0$ corresponds to Poissonian statistics typical of coherent states, $\delta n^{\rm ss}_{\ell}>0$ corresponds to super-Poissonian statistics, and sub-Poissonian statistics with $\delta n^{\rm ss}_{\ell}<0$ are known to be incompatible with a classical description of the state [61].
Focusing on the last site, $\ell=L$, the results are collected in Fig. 3. Qualitatively similar results are obtained throughout the chain, and we refer the reader to the Supplementary Information for results at $\ell=1$ and $\ell=L/2$. We also provide representative instances of local Wigner functions $W_{L}^{\rm ss}(\alpha,\alpha^{*})$. We identify three distinct regimes:
Figure 4: Spatial relaxation of amplitude and phase degrees of freedom in the chaotic regime. Spatial profiles of equal-time statistics of the complex field, decomposed as $\alpha_{\ell}=:r_{\ell}\,{\rm e}^{{\rm i}\phi_{\ell}}$ with $\ell=1,...L$, once the chaotic NESS is reached.
(a) Average amplitude $\langle r_{\ell}\rangle_{\rm ss}$, showing a large and increasing relaxation scale for increasing chain lengths $L=$ 80, 200, 300, 400.
(b) Average phase $\langle\phi_{\ell}\rangle_{\rm ss}$, rapidly converging to zero. (c) First-order coherence function $|g_{k,\ell}^{(1)}|$ defined in Eq. (7) for $k=L$ (blue) and $k=L/2$ (light-blue), showing sharp exponential decorrelation of the phases away from $\ell=k$. The drive strength is set to $F=7.5$ and the other parameters are set as in Fig. 2.
  1. (I)
    Regular quasilinear regime; at weak drive, only few photons populate the last site. The photon statistics are Poissonian ($\delta n^{\rm ss}_{L}\approx 0$) and the small saturation value of the phase OTOC, $D^{\rm ss}_{1,L}(\tau\to\infty)\approx 10^{-2}-10^{-1}$, indicates regular dynamics. The Wigner function exhibits the Gaussian shape characteristic of a coherent state. In this regime, single-particle excitations are dilute, rendering nonlinearities negligible and preventing the onset of chaos.
  2. (II)
    Chaotic regime; at stronger drives, the photon number markedly increases with respect to the quasilinear regime and nonlinearities are now relevant. The large value of $\delta n^{\rm ss}_{L}$ indicates strong super-Poissonian fluctuations. Concurrently, the phase OTOC saturates to $D^{\rm ss}_{1,L}(\tau\to\infty)\approx 1$, signaling significant dephasing between the fields in the two replicas. This is also reflected in the local Wigner function where the ergodicity of Wigner trajectories translates into $W_{L}^{\rm ss}(\alpha,\alpha^{*})$ spreading across a wide portion of phase space. These observations, along with the exponential growth of $D_{1,\,\ell}^{\rm ss}(\tau)$ shown in Fig. 2 (a) for a representative value of $F$ indicated by a pentagon marker, confirm the chaotic nature of the quantum dynamics.
  3. (III)
    Regular coherent regime; at even stronger drives, the photon number statistics become sub-Poissonian ($\delta n_{L}<0$). This regime is incompressible in the sense that the photon number marginally increases with $F$ [62]. The Wigner function acquires a crescent-shaped distortion with respect to the coherent state [44] which is a typical signature of many-body phases driven by the competition of Kerr-like nonlinearities and detuning, such as in optical bistability [63]. The small saturation value $D_{1,\,L}^{\rm ss}(\tau\to\infty)\approx 10^{-2}-10^{-1}$ indicates persistent phase correlations between the two replicas. These observations, along with the sub-exponential growth of $D_{1,\,\ell}^{\rm ss}(\tau)$ shown in Fig. 2 (b) for a representative value of $F$ indicated by a square marker, are indicative of regular dynamics.
Notably, the crossover from the regular quasilinear regime (I) to the chaotic regime (II) is smooth in both $n^{\rm ss}_{L}$ and $\delta n^{\rm ss}_{L}$ while the transition between the chaotic regime (II) and the regular coherent regime (III) is characterized by abrupt variations. In particular, the drive strength $F$ at which the photon number leaps coincides with that of the maximum of $\delta n^{\rm ss}_{L}$.
II.3 Two-stage thermalization in space
In this section, we focus on the chaotic regime by fixing the drive strength to $F=7.5$. We decompose the complex field in terms of amplitude and phase degrees of freedom, $\alpha_{\ell}=r_{\ell}\,{\rm e}^{{\rm i}\phi_{\ell}}$ where $r_{\ell}\geq 0$, and we separately analyze their spatial relaxation along the chain, from $\ell=1$ to $\ell=L$.
Figure 4 shows the steady-state profile of the average amplitude $\langle r_{\ell}\rangle_{\rm ss}$ and average phase $\langle\phi_{\ell}\rangle_{\rm ss}$ across the chain for various chain lengths $L$. The amplitude of the field slowly decays along the chain. This relaxation is characterized by a length scale $\xi_{r}$ that clearly grows monotonically with the system size $L$. This large length scale can be explained by the local conservation of the photon number within the bulk of the chain, which hinders the rapid relaxation of the amplitude degree of freedom. Instead, this relaxation occurs through much slower hydrodynamic processes involving, notably, the driven-dissipative conditions at the two boundaries of the chain. In contrast, the phase degree of freedom averages to zero over a much shorter, microscopic, length scale $\xi_{\phi}$ spanning only a few sites.
To corroborate the short relaxation of the phase sector, we compute the spatial correlations of the phases in the chain by means of the first-order coherence function,
$g^{(1)}_{k,\ell}:=\frac{\langle\hat{a}_{k}^{\dagger}\hat{a}_{\ell}\rangle_{\rm ss}}{\sqrt{\langle\hat{a}_{k}^{\dagger}\hat{a}_{k}\rangle_{\rm ss}\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle_{\rm ss}}}\,.$ (7)
The results are presented in Fig. 4 (c) as a function of $\ell$, for both $k=L/2$ and $k=L$. In both cases, the profile of the spatial correlations shows exponential decay, $|g^{(1)}_{k,\ell}|\sim{\rm e}^{-|k-\ell|/\xi_{\phi}}$, typical of a disordered phase with a correlation length $\xi_{\phi}$ on the order of a few sites, before saturating to values below $10^{-2}$.
Altogether, these results suggest a two-stage thermalization process in space: the phase sector thermalizes to a disordered state over a short length scale, while the amplitude sector is contingent upon both boundary conditions and relaxes over the entire system. As we shall demonstrate below, this opens up a significant region of space, beyond the first few sites near the boundary drive, but before the complete relaxation to near vacuum states subject to thermal fluctuations, where the U(1) symmetry of the undriven Hamiltonian is restored and the highly-populated sites along the chain can be seen as in local equilibria characterized by a large and slowly-varying chemical potential.
Figure 5: Prethermal phase in the chaotic regime. (a) Spatial profile of the quantity $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}$ across a chain of length $L=400$. $\beta^{\rm eff}_{\ell}$ and $\mu^{\rm eff}_{\ell}$ are the effective local inverse temperature and chemical potential, respectively. The results are obtained by mapping the local physics at site $\ell$ to the 2-photon impurity ansatz defined in Eqs. (8-9) and the quantity $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}$ is computed via Eq. (10). We also report the results obtained by mapping to the generalized Scully-Lamb ansatz and to an equilibrium Gibbs measure. Three distinct domains are discussed in the text: 1) nonsymmetric nonthermal domain where the ansatz fails to reproduce $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$ such as the one in panel (c) ; 2) prethermal domain with $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}>0$ ; 3) thermal domain with $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}\leq 0$. (b) Respective sizes of the three domains versus the overall chain length $L$. (c-e) Normalized local Wigner functions, $\tilde{W}_{\ell}^{\rm ss}(\alpha,\alpha^{*}):=W^{\rm ss}_{\ell}(\alpha,\alpha^{*})/\textrm{max}[W^{\rm ss}_{\ell}(\alpha,\alpha^{*})]$ for representative sites in the three domains of a chain of length $L=400$ and corresponding to the markers in panel (a). The drive strength is set to $F=7.5$ and the other parameters are set as in Fig. 3.
We test this scenario by probing the local symmetries, the thermodynamics, and the temporal fluctuations across the chain. On general grounds, the local steady-state density matrix $\hat{\rho}^{\rm ss}_{\ell}:=\lim\limits_{t\to\infty}\textrm{Tr}_{k\neq\ell}\hat{\rho}(t)$ can be cast as the solution of an effective driven-dissipative impurity model constructed by singling out one site of the chain, and tracing out its neighbors such as to treat them as incoherent sources and drains. In perfect thermal equilibrium with U(1) symmetry, the statics of the impurity model are to be straightforwardly described by the Gibbs state $\hat{\rho}^{\rm eq}_{\ell}\propto\exp\small[-\beta(\hat{H}_{\rm imp}-\mu\,\hat{a}^{\dagger}a)]$ where $\beta$ and $\mu$ are the inverse temperature and chemical potential, respectively, and the impurity Hamiltonian
$\hat{H}_{\rm imp}=\omega_{0}\,\hat{a}^{\dagger}\hat{a}+\frac{1}{2}U\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}\,,$ (8)
corresponds to a single site of the chain Hamiltonian in Eq. (2) (see the Supplementary Information). In a generic NESS, however, the rigorous derivation of an effective driven-dissipative impurity model is a daunting task, and we simply rely on symmetries and heuristics to identify a minimal impurity model that successfully locally captures both the statics as well as temporal fluctuations. Assuming that the U(1) symmetry is indeed rapidly restored in the chain, this naturally brings us to consider the impurity ansatz defined by the local U(1)-symmetric Hamiltonian $\hat{H}_{\rm imp}$ in Eq. (8) and by the first non-trivial jump operators allowed under weak U(1) symmetry [64]
$\hat{L}_{\uparrow}\!=\!\sqrt{\gamma^{\uparrow}}\hat{a}^{\dagger}\!,\hat{L}_{\downarrow}\!=\!\sqrt{\gamma^{\downarrow}}\hat{a},\hat{L}_{\phi}\!=\!\sqrt{\gamma^{\phi}}\hat{a}^{\dagger}\hat{a},\hat{L}_{\rm s}\!=\!\sqrt{\gamma^{\rm s}}\hat{a}^{2}.$ (9)
The non-negatives parameters $\gamma^{\uparrow}$, $\gamma^{\downarrow}$, $\gamma^{\phi}$, and $\gamma^{\rm s}$ are effective rates of incoherent pumping, decay, dephasing, and 2-photon decay. The inclusion of the latter is essential to ensure the saturation of the photon number when $\gamma^{\uparrow}>\gamma^{\downarrow}$.
We test the validity of this effective modeling, and the underlying restoration of the U(1) symmetry, by fitting the local steady-state Wigner function $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$ along the chain to the ones predicted by the impurity ansatz. The fitting parameters are $\gamma^{\uparrow}$, $\gamma^{\downarrow}$ and $\gamma^{\rm s}$ since the steady-state Wigner function of the impurity ansatz is independent of the Hamiltonian parameters and of the dephasing rate $\gamma^{\phi}$. The latter can be extracted by fitting two-time correlation functions. The overall fitting procedure is detailed in the Supplementary Information. We repeat this procedure at all sites from $\ell=1$ to $\ell=L$. Excellent matches are obtained everywhere along the chain except for the first few sites, confirming the scenario of the U(1) symmetry being rapidly restored away from the boundary drive.
Let us stress that the above steady-state impurity modeling is not unique. Notably, we also tested a generalized version of the Scully-Lamb model [65] defined by $\hat{L}_{\rm s}=0$ and a modified $\hat{L}_{\uparrow}={\hat{a}^{\dagger}(\gamma^{\uparrow}-\mathcal{S}\,\hat{a}\hat{a}^{\dagger})}/{\sqrt{\gamma^{\uparrow}}}$, where $\mathcal{S}\geq 0$ is a photon saturation rate. In the context of lasing, this model can be derived as an effective theory for the optical degree of freedom when the (inverted) atomic population modeled by two-level systems has been integrated out [66, 67]. This ansatz proved to be equally successful in capturing $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$, yielding comparable values of the effective parameters $\gamma^{\uparrow}$ and $\gamma^{\downarrow}$.
II.4 Prethermal phase
We now use the above mapping to the driven-dissipative impurity ansatz defined in Eqs. (8-9) as a local thermometer along the chain. The effective inverse temperature $\beta^{\rm eff}_{\ell}$ and chemical potential $\mu^{\rm eff}_{\ell}$ at site $\ell$ can be related to the single-photon rates via the detailed-balance condition (we set $k_{\rm B}=1$)
$\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}:=\log(\gamma^{\uparrow}_{\ell}\big{/}\gamma^{\downarrow}_{\ell})\,,$ (10)
where $\gamma^{\downarrow}_{\ell}$ and $\gamma^{\uparrow}_{\ell}$ are the parameters of the impurity model determined at site $\ell$. We absorbed all the quadratic contributions to $\hat{H}_{\rm imp}$, namely $\omega_{0}-U/2$, into a redefinition of the effective chemical potential.
In Fig. 5 (a), we plot the product $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}$ as a function of $\ell$ and we identify three distinct spatial domains:
  1. 1.
    Nonsymmetric nonthermal domain; as discussed above, the first few sites are not captured by the impurity ansatz because of the proximity to the U(1)-breaking drive. This breaking of U(1) is visible in the asymmetry of local Wigner functions represented in Fig. 5 (c). Here, the quantity $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}$ is ill-defined, as indicated by the shaded region in the figure.
  2. 2.
    U(1)-symmetric prethermal domain; in this intermediate region of the chain, local quantities are well captured by the impurity ansatz with a positive chemical potential, $\mu^{\rm eff}_{\ell}>0$, and a finite 2-photon decay rate $\gamma^{\rm s}$. The relatively constant number of photons reported in Fig. 4 (a), along with the large ring-shaped Wigner functions in Fig. 5 (d), can be interpreted as the result of the competition between a large chemical potential that strives to add photons and the Kerr non-linearity that acts as a saturation mechanism. However, this state should not be identified to lasing since here, unlike lasers, the temporal phase coherence is short-lived, indicating that the U(1) symmetry breaking characteristic of lasing has not occurred (see the Supplementary Information). Instead, the phase evolves diffusively [68], restoring the U(1) symmetry of the chain Hamiltonian.
  3. 3.
    U(1)-symmetric thermal domain; the right side of the chain is captured by the impurity ansatz with a negative chemical potential, $\mu^{\rm eff}_{\ell}\leq 0$. The corresponding Wigner functions, illustrated in Fig. 5 (e), display bell-shaped envelops that can be interpreted as the outcome of a competition between the chemical potential depleting the photon population towards the vacuum, and the thermal fluctuations that sustain a residual population of weakly-interacting photons.
To test the robustness of this hydrodynamic description, we repeated the analysis above using the generalized Scully-Lamb ansatz instead of the 2-photon ansatz, as well as directly fitting the local Wigner functions with those of Gibbs states
$\hat{\rho}^{\rm eq}_{\ell}=\frac{\exp[-\beta_{\ell}^{\rm eff}(\hat{H}_{\rm imp}-\mu_{\ell}\,\hat{a}^{\dagger}a)]}{\operatorname{Tr}\left(\exp[-\beta_{\ell}^{\rm eff}(\hat{H}_{\rm imp}-\mu_{\ell}\,\hat{a}^{\dagger}a)]\right)}\,.$ (11)
A detailed comparison between these three thermometers can be found in the Supplementary Information. As reported in Fig. 5 (a), we obtained an excellent qualitative agreement between these approaches and even quantitative agreement in the thermal phase.
Remarkably, we argue in Fig. 5 (b) that the sizes of both the nonsymmetric and thermal domains are limited to sub-dominant portions of the chain, whereas the size of the prethermal domain increases linearly with the overall chain length $L$. This suggests that the latter may be an extensive thermodynamic phase that occurs ahead of the eventual complete thermalization at the rightmost portion of the chain.
III Discussion
The NESS of the boundary-driven dissipative chain of nonlinear oscillators revealed an extensive and chaotic region of the chain that can be effectively described hydrodynamically in terms of local equilibria with a large, emergent, chemical potential. Drawing an analogy with the long-lived prethermal phases observed in the temporal relaxation of many-body systems following a quench, this can be interpreted as a prethermal phase of matter that builds up in space rather than in time.
An immediate question concerns how this prethermal phase influences the transport properties of the driven-dissipative Bose-Hubbard chain [31, 32, 69, 70, 71]. Akin to the thermodynamic description obtained with the single-site ansatz, transport properties could be captured and analyzed using multi-site impurity models. It remains to be determined whether the length of the prethermal domain scales thermodynamically, or instead saturates at a large but finite value in very large chains. The former scenario seems plausible if the slow relaxation of the steady-state photon density across the chain is primarily governed by the distant driven-dissipative conditions at the two boundaries of the chain. However, the latter scenario is supported by the analogy to standard prethermalization mechanisms, where the timescale for escaping the prethermal state typically follows an exponential dependence, $\tau\sim\exp(1/\delta)$, with $\delta$ quantifying the strength of the integrability-breaking perturbation—here, characterized by $U/J$. Resolving this question, along with assessing the relevance of next-to-leading-order quantum fluctuations that were neglected so far, requires studying system sizes $L\gg\exp(J/U)$, well beyond current experimental capabilities. This challenge will necessitate significantly larger numerical simulations or the development of new methodologies that extend beyond the TWA to effectively handle intermediate Kerr nonlinearities while still incorporating quantum fluctuations.
More broadly, our proposed mechanism indicates that the emergence of a prethermal chaotic phase hinges on three essential conditions: first, an interacting bulk with at least one conserved charge and a sufficiently large and complex local Hilbert space that can facilitate the relaxation of different degrees of freedom across well-separated length scales; second, a local nonequilibrium drive that explicitly breaks the symmetry associated with the conservation law; and finally, weakly-symmetric dissipation channels positioned far from the drive, ensuring nontrivial steady states and a prolonged hydrodynamic relaxation of the conserved charge. These minimal ingredients, along with the semiclassical methodology we introduced, lay the foundation for identifying prethermal chaos in a wide range of boundary-driven systems.
IV Methods
IV.1 Stochastic semiclassical description
The TWA is a semiclassical approximation of the quantum many-body dynamics that accounts for leading-order quantum fluctuations. It relies on a phase-space representation of the system’s density matrix $\hat{\rho}$ in terms of the Wigner function $W(\alpha_{1},\alpha_{1}^{*},...,\alpha_{L},\alpha_{L}^{*})$, where $\alpha_{\ell}$ and $\alpha_{\ell}^{*}$ for $\ell=1,...,L$ are the complex amplitudes associated with the local coherent states. In this framework, the Lindblad master equation on the operator $\hat{\rho}$ in Eq. (1) is mapped to a partial differential equation on $W$. Notably, two-body interactions in the Hamiltonian yield contributions up to third order derivatives of the type $\alpha_{\ell}\,\partial^{3}W/\partial\alpha^{*}_{\ell}\partial^{2}\alpha_{\ell}$. The approximation consists in discarding third and higher-order derivatives, reducing the dynamics to a Fokker-Plank equation where $W$ can be interpreted as a probability distribution of the phase-space variables. The latter equation can then be unraveled into a set of $L$ coupled stochastic differential equations on the complex amplitudes $\alpha_{\ell}$ given by Eq. (II.1). In practice, we compute the solutions of these Langevin-like equations, the so-called Wigner trajectories, by means of numerical solvers specific to stochastic differential equations. Observables are computed by averaging over a large number of trajectories generated by different realizations of the quantum noise. The dictionary between the original Lindblad master equation framework and the Wigner-trajectory implementation of the TWA framework reads, notably,
Observable Lindblad Wigner trajectories
Field $\langle\hat{a}_{\ell}\rangle$ $\operatorname{Tr}[\hat{a}_{\ell}\hat{\rho}]$ $\langle\alpha_{\ell}\rangle$
Photon number $\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle$ $\operatorname{Tr}[\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\hat{\rho}]$ $\langle|\alpha_{\ell}|^{2}\rangle-1/2$
Spatial correlation $\langle\hat{a}_{k}^{\dagger}\hat{a}_{\ell}\rangle$ $\operatorname{Tr}[\hat{a}_{k}^{\dagger}\hat{a}_{\ell}\hat{\rho}]$ $\langle\alpha_{k}^{*}\alpha_{\ell}\rangle-\delta_{kl}/2$
Kerr nonlinearity $\langle\hat{a}_{\ell}^{\dagger 2}\hat{a}_{\ell}^{2}\rangle$ $\operatorname{Tr}[\hat{a}_{\ell}^{\dagger 2}\hat{a}_{\ell}^{2}\hat{\rho}]$ $\langle|\alpha_{\ell}|^{4}\rangle-2\langle|\alpha_{\ell}|^{2}\rangle+1/2$
where delimited-⟨$\langle$..$\rangle$ … ⟩ in the observable column is the standard quantum expectation value and, in the Wigner trajectories column, denotes the average with respect to Wigner trajectories. The local Wigner function $W_{\ell}(t;\alpha,\alpha^{*})$ is a phase-space representation of the reduced density matrix at site $\ell$, $\hat{\rho}_{\ell}(t):=\textrm{Tr}_{k\neq\ell}\,\hat{\rho}(t)$. It can be simply reconstructed by generating the histogram of $\alpha_{\ell}(t)$ in the complex plane when sampling over Wigner trajectories. When the NESS is reached, single-time observables converge to constant values and $W_{\ell}(t;\alpha,\alpha^{*})\to W^{\rm ss}_{\ell}(\alpha,\alpha^{*})$. There, the statistics can be improved by also sampling the trajectories in time, significantly reducing the computational overhead.
Figure 6: Scrambling of information in the NESS. Space-time evolution of the steady-state phase OTOC $D_{1,\ell}^{\rm ss}(\tau)$, defined in Eq. (14) and computed via Eq. (5), across an $L=80$ chain in the chaotic regime. The white dashed line indicates ballistic spreading of information at the butterfly velocity $v=2J$. The results are obtained upon averaging over $5\times 10^{3}$ independent Wigner trajectories. The drive strength has been fixed to $F=7.5$ and the other parameters are set as in Fig. 2.
IV.2 Phase OTOC
Here, we justify the expression of the phase OTOC $D_{k,\ell}(t,\tau)$ given in Eq. (5) by deriving it as the semiclassical limit of the following quantum OTOC,
$C_{k,\ell}(t,\tau)\!:=\!\frac{1}{2}\textrm{Tr}\Big{(}[\hat{W}_{\ell}(t\!+\!\tau),\hat{V}_{k}(t)]^{\dagger}[\hat{W}_{\ell}(t\!+\!\tau),\hat{V}_{k}(t)]\hat{\rho}(t)\Big{)}.$ (12)
[…]
$\{f,g\}:=\sum_{\ell=1}^{L}\left(\frac{\partial f}{\partial n_{\ell}}\frac{\partial g}{\partial\phi_{\ell}}-\frac{\partial g}{\partial n_{\ell}}\frac{\partial f}{\partial\phi_{\ell}}\right)\,.$ (13)
Carrying out this replacement in Eq. (12), using the relation $\{\phi_{\ell}(t^{\prime}),n_{k}(t)\}=-\delta\phi_{\ell}(t^{\prime})/\delta\phi_{k}(t)$, one obtains the semiclassical version of $C_{k,\ell}(t,\tau)$ which reads
$D_{k,\ell}(t,\tau)$ $=\frac{1}{2}\left\langle\left|\frac{\delta{\rm e}^{{\rm i}\phi_{\ell}(t+\tau)}}{\delta\phi_{k}(t)}\right|^{2}\right\rangle\,,$ (14)
where delimited-⟨$\langle$..$\rangle$ … ⟩ denote the average over the quantum noise and ${\delta}/{\delta\phi_{k}(t)}$ implements an infinitesimal perturbation of the phase at site $k$ and time $t$. In practice, this is implemented by cloning the system in two replicas $a$ and $b$, applying an infinitesimal perturbation at time $t$ on the phase at site $k$ of replica $b$, subsequently evolving both replicas subject to the same quantum noise, and finally averaging over realizations of the quantum noise. Thus, the semiclassical phase OTOC can be cast as
$D_{k,\ell}(t,\tau)=\frac{1}{2}\left\langle\left|{\rm e}^{{\rm i}\phi_{\ell}^{(a)}(t+\tau)}-{\rm e}^{{\rm i}\phi_{\ell}^{(b)}(t+\tau)}\right|^{2}\right\rangle\,,$ (15)
which boils down to the operational definition given in Eq. (5).
Whenever the dynamics reaches a steady state, $\hat{\rho}_{\rm ss}:=\lim\limits_{t\to\infty}\hat{\rho}(t)$, we can define the semiclassical steady-state phase OTOC as $D^{\rm ss}_{k,\ell}(\tau):=\lim\limits_{t\to\infty}D_{k,\ell}(t,\tau)$. As a sanity check to back up the use of $D^{\rm ss}_{k,\ell}(\tau)$ as a dynamical probe of quantum chaos in the NESS, we verify that it reproduces the ballistic spreading of information one expects of a tight-binding model. Figure 6 clearly exhibits a causal light-cone structure with a finite butterfly velocity $v=2J$.
Acknowledgements.
We acknowledge enlightening discussions with Alberto Biella, Andreas Läuchli, Alberto Mercurio, Luca Gravina, Lorenzo Fioroni and Pasquale Scarlino. CA acknowledges the support from the French ANR “MoMA” project ANR-19-CE30-0020 and the support from the project 6004-1 of the Indo-French Centre for the Promotion of Advanced Research (IFCPAR). FF, FM, and VS acknowledge the support by the Swiss National Science Foundation through Projects No. 200020_185015, 200020_215172, and 20QU-1_215928, and by the EPFL Science Seed Fund 2021.
Supplementary information
1. Numerical approaches
Quantum trajectories
Our starting point is the Lindblad master equation written in its most general form
$\frac{\partial\hat{\rho}}{\partial t}=-{\rm i}[\hat{H},\hat{\rho}]+\sum_{j}\left(\hat{L}_{j}\hat{\rho}\hat{L}_{j}^{\dagger}-\frac{1}{2}\left\{\hat{L}_{j}^{\dagger}\hat{L}_{j},\hat{\rho}\right\}\right),$ (16)
where $\{\hat{L}_{j}\}$ is a collection of Lindblad jump operators. Equation (16) admits a stochastic unraveling in terms of quantum trajectories |$\ket{\psi(t)}$, combining the Hamiltonian dynamics with a continuous monitoring of the environment [72, 73, 74]. Different measurement’s protocols (i.e., different unraveling protocols) are possible, and they lead to different quantum trajectories. The most popular protocols are the homodyne measurement, which results in a stochastic Wiener process for the system’s state |$\ket{\psi(t)}$ [75], and the photon-counting measurement, giving rise to the Monte Carlo quantum trajectories [76]. In the latter case, a quantum jump occurs in a time step ${\rm d}t$ with probability ${\rm d}p=\sum_{j}\bra{\psi(t)}\hat{L}_{j}^{\dagger}\hat{L}_{j}\ket{\psi(t)}{\rm d}t$, and |$\ket{\psi(t)}$ evolves into
|$\ket{\psi(t+{\rm d}t)}\propto\hat{L}_{j}\ket{\psi(t)},$ (17)
where the jump operator $\hat{L}_{j}$ is sampled from the probability distribution
$p_{j}=\frac{\bra{\psi(t)}\hat{L}_{j}^{\dagger}\hat{L}_{j}\ket{\psi(t)}}{\sum_{k}\bra{\psi(t)}\hat{L}_{k}^{\dagger}\hat{L}_{k}\ket{\psi(t)}}.$ (18)
No quantum jump occurs in the time ${\rm d}t$ with probability $1-{\rm d}p$, and |$\ket{\psi(t)}$ evolves into
|$\ket{\psi(t+{\rm d}t)}\propto(\hat{1}-{\rm i}\,{\rm d}t\hat{H}_{\rm nh})\ket{\psi(t)},$ (19)
where $\hat{H}_{\rm nh}:=\hat{H}-{\rm i}\sum_{j}\hat{L}_{j}^{\dagger}\hat{L}_{j}/2$ is the associated non-Hermitian Hamiltonian. After each time step, the state is renormalized according to |$\ket{\psi(t+{\rm d}t)}\mapsto\ket{\psi(t+{\rm d}t)}/\langle\psi(t+{\rm d}t)|\psi(t+{\rm d}t)\rangle$.
The overall process leads to a stochastic Schrödinger equation for the wave function |$\ket{\psi(t)}$. Expectation values of operators can be obtained by averaging over many independent quantum trajectories (see the discussion below).
Truncated Wigner approximation
In what follows, we mainly adapt the discussion on phase-space representation presented in Ref. [39]. To simplify the discussion, we initially consider a single-mode resonator described by a Hamiltonian $\hat{H}$ and a collection of Lindblad jump operators $\hat{L}_{j}$ that can be written as polynomials of single-mode creation and annihilation operators $\hat{a}^{\dagger}$ and $\hat{a}$. The phase-space representation maps operators (defined in an $M$-dimensional Hilbert space) into functions (defined in the phase space). Here, we work in the coherent-state basis {|$\{\ket{\alpha}\}$, where |$\ket{\alpha}$ is the eigenstate of the annihilation operator $\hat{a}$ with eigenvalue $\alpha$. The one-to-one mapping between an operator $\hat{O}$ in the Hilbert space and a function $O_{W}$ in phase space can be achieved by introducing the Weyl symbol
$O_{W}(\alpha,\alpha^{*}):=\frac{1}{2^{M}}\int{\rm d}\zeta{\rm d}\zeta^{*}\bra{\alpha-\frac{1}{2}\zeta}\hat{O}(\hat{a},\hat{a}^{\dagger})\ket{\alpha+\frac{1}{2}\zeta}{\rm e}^{-\left(\alpha^{*}-\frac{\zeta^{*}}{2}\right)\left(\alpha+\frac{\zeta}{2}\right)}.$ (20)
If the operator $\hat{O}(\hat{a},\hat{a}^{\dagger})$ is symmetric in $\hat{a}$ and $\hat{a}^{\dagger}$, then the Weyl symbol can be obtained with the simple substitution $\hat{a}\rightarrow\alpha$, $\hat{a}^{\dagger}\rightarrow\alpha^{*}$. The Weyl symbol of the density matrix $\hat{\rho}$ is called the Wigner function,
$W(\alpha,\alpha^{*}):=\int\frac{{\rm d}\zeta{\rm d}\zeta^{*}}{2\pi}\bra{\alpha-\frac{1}{2}\zeta}\hat{\rho}\ket{\alpha+\frac{1}{2}\zeta}{\rm e}^{-\left(\alpha^{*}-\frac{\zeta^{*}}{2}\right)\left(\alpha+\frac{\zeta}{2}\right)}.$ (21)
Within this formalism, expectation values of operators can be expressed as
$\langle\hat{O}(\hat{a},\hat{a}^{\dagger})\rangle=\int{\rm d}\alpha{\rm d}\alpha^{*}\,O_{W}(\alpha,\alpha^{*})W(\alpha,\alpha^{*}),$ (22)
i.e., classical statistical expectation values weighted over the Wigner function. The Weyl symbol corresponding to a product of operators can be expressed as
$(\hat{O}_{1}\hat{O}_{2})_{W}(\alpha,\alpha^{*})=O_{1W}(\alpha,\alpha^{*})\,{\rm e}^{\overset{\leftrightarrow}{\Lambda}/2}\,O_{2W}(\alpha,\alpha^{*}),$ (23)
where ${\rm e}^{\overset{\leftrightarrow}{\Lambda}/2}$ is a Moyal product based on the derivative operator associated with the Poisson bracket
Λ:=
α
α
α
α
assignΛ
[…]
$([\hat{O}_{1},\hat{O}_{2}])_{W}(\alpha,\alpha^{*})=2\,O_{1W}(\alpha,\alpha^{*})\sinh\left(\frac{1}{2}\overset{\leftrightarrow}{\Lambda}\right)O_{2W}(\alpha,\alpha^{*}).$ (25)
Weyl symbols of any operator $\hat{O}(\hat{a},\hat{a}^{\dagger})$ can be computed using the Moyal product in Eq. (23). For example,
(a^a^)W=α(112
α
α
α
=|α|212
[…]
(a^a^a^a^)W=α2(112
α
α
+18
2
α2
2α2
α2
=|α|42|α|2+12
[…]
18superscript
2
[…]
Let us notice that the zeroth-order expansion of the Moyal product would have led to $(\hat{a}^{\dagger}\hat{a})_{W}=|\alpha|^{2}$ and $(\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a})_{W}=|\alpha|^{4}$, i.e., treating quantum operators classical commuting quantities. The remaining terms in Eqs. (26) and (27) originate from the quantum fluctuations that are captured by the systematic expansion of the Moyal product.
The phase-space representation can be easily generalized to spatially extended systems, where the full Hilbert space is now the tensor product of, say, $L$ local Hilbert spaces. The Wigner function associated with the density matrix $\hat{\rho}(t)$ is now defined as
$W(t;\alpha_{1},\alpha_{1}^{*},...,\alpha_{L},\alpha_{L}^{*}):=$ $\int\frac{{\rm d}\zeta_{1}{\rm d}\zeta_{1}^{*}...{\rm d}\zeta_{L}{\rm d}\zeta_{L}^{*}}{(2\pi)^{L}}\big{\langle}\alpha_{1}-\frac{\zeta_{1}}{2},...,\alpha_{L}-\frac{\zeta_{L}}{2}\big{|}\hat{\rho}(t)\big{|}\alpha_{1}+\frac{\zeta_{1}}{2},...,\alpha_{L}+\frac{\zeta_{L}}{2}\big{\rangle}$
×$\quad\times{\rm e}^{-\left(\alpha_{1}^{*}-\frac{1}{2}\zeta_{1}^{*}\right)\left(\alpha_{1}+\frac{1}{2}\zeta_{1}\right)}...\,{\rm e}^{-\left(\alpha_{L}^{*}-\frac{1}{2}\zeta_{L}^{*}\right)\left(\alpha_{L}+\frac{1}{2}\zeta_{L}\right)}.$ (28)
Similarly, the expressions for the Weyl symbols and expectation values are easily generalized from the single-mode case in Eqs. (20) and (22). Finally, the Poisson bracket in Eq. (24) generalizes to
Λ==1L(
α
α
α
α
[…]
Within this theoretical framework, the Lindblad equation (1) can be mapped into a partial differential equation (PDE) for the Wigner function $W(t;\alpha_{1},\alpha^{*}_{1},...,\alpha_{L},\alpha^{*}_{L})$ that reads
${\rm i}\frac{\partial W}{\partial t}=$ $-\Delta\sum_{\ell=1}^{L}\left(\alpha_{\ell}^{*}\frac{\partial}{\partial\alpha_{\ell}^{*}}-\alpha_{\ell}\frac{\partial}{\partial\alpha_{\ell}}\right)W-J\sum_{\ell=1}^{L-1}\left(\alpha_{\ell+1}^{*}\frac{\partial}{\partial\alpha_{\ell}^{*}}-\alpha_{\ell+1}\frac{\partial}{\partial\alpha_{\ell}}\right)W-F\left(\frac{\partial}{\partial\alpha_{1}^{*}}-\frac{\partial}{\partial\alpha_{1}}\right)W$
$+U\sum_{\ell=1}^{L}(|\alpha_{\ell}|^{2}-1)\left(\alpha_{\ell}^{*}\frac{\partial}{\partial\alpha_{\ell}^{*}}-\alpha_{\ell}\frac{\partial}{\partial\alpha_{\ell}}\right)W-\frac{U}{4}\sum_{\ell=1}^{L}\left(\alpha_{\ell}^{*}\frac{\partial}{\partial\alpha_{\ell}^{*}}-\alpha_{\ell}\frac{\partial}{\partial\alpha_{\ell}}\right)\frac{\partial^{2}}{\partial\alpha_{\ell}\partial\alpha_{\ell}^{*}}W$
$+\frac{{\rm i}\gamma}{2}\left[\frac{\partial}{\partial\alpha_{1}}(\alpha_{1}W)+\frac{\partial}{\partial\alpha_{1}^{*}}(\alpha_{1}^{*}W)+\frac{\partial^{2}}{\partial\alpha_{1}\partial\alpha_{1}^{*}}W+\frac{\partial}{\partial\alpha_{L}}(\alpha_{L}W)+\frac{\partial}{\partial\alpha_{L}^{*}}(\alpha_{L}^{*}W)+\frac{\partial^{2}}{\partial\alpha_{L}\partial\alpha_{L}^{*}}W\right].$ (30)
The mapping of Eq. (1) to Eq. (Truncated Wigner approximation) is exact. The TWA consists in expanding the Moyal product ${\rm e}^{\overset{\leftrightarrow}{\Lambda}/2}$ up to second order and neglecting higher-order quantum fluctuations. In Eq. (Truncated Wigner approximation), this amounts to discarding the third-order derivative terms that stem from the Kerr nonlinearity while keeping lower-order contributions. This semiclassical approximation is valid in the limit of weak Kerr nonlinearity [40]. Within the TWA, the PDE becomes a Fokker-Planck equation [38] and $W$ can be interpreted as a well-defined probability distribution of the phase space variables. Finally, the Fokker-Planck equation can be mapped to a Langevin equation [38], yielding the set of stochastic differential equations (II.1). Expectation values of operators can be mapped from integrals over the Wigner function (22) into statistical expectation values of the Weyl symbols averaged over many Wigner trajectories solutions of Eqs. (II.1),
$\int\!{\rm d}\alpha_{1}{\rm d}\alpha^{*}_{1}...{\rm d}\alpha_{L}{\rm d}\alpha^{*}_{L}O_{W}(\alpha_{1},\alpha^{*}_{1},...,\alpha_{L},\alpha^{*}_{L})W(t;\alpha_{1},\alpha^{*}_{1},...,\alpha_{L},\alpha^{*}_{L})=\frac{1}{N_{\textrm{traj}}}\!\sum_{j=1}^{N_{\textrm{traj}}}O_{W}\big{(}\alpha_{1}^{(j)}(t),\alpha_{1}^{*(j)}(t),...,\alpha_{L}^{(j)}(t),\alpha_{L}^{*(j)}(t)\big{)}.$ (31)
Finally, the Wigner trajectories also give access to the local Wigner function $W_{\ell}(t;\alpha,\alpha^{*})$ that is defined as
$W_{\ell}(t;\alpha_{\ell},\alpha^{*}_{\ell}):=\int\!{\rm d}\alpha_{1}\alpha_{1}^{*}...{\rm d}\alpha_{\ell-1}{\rm d}\alpha_{\ell-1}^{*}{\rm d}\alpha_{\ell+1}{\rm d}\alpha_{\ell+1}^{*}...{\rm d}\alpha_{L}{\rm d}\alpha^{*}_{L}\,W(t;\alpha_{1},\alpha^{*}_{1},...,\alpha_{L},\alpha^{*}_{L}).$ (32)
To construct the steady-state quantity $W^{\rm ss}_{\ell}(\alpha,\alpha^{*}):=\lim\limits_{t\to\infty}W_{\ell}(t;\alpha,\alpha^{*})$, we realize a histogram from the data of $5\times 10^{7}$ fields $\alpha_{\ell}(t)$. The fields have been obtained from the time evolution of $20$ Wigner trajectories after average quantities have reached steady-state values, sampling over a time window $\Delta\tau=2\times 10^{5}$. Those numbers are fixed for all the Wigner functions presented in the manuscript.
The numerical results have been obtained by numerically solving Eqs. (II.1) using the solver SOSRI of the package Stochastic Differential Equations that is available in Julia. The initial conditions $\alpha_{\ell}(0)$ for each trajectory are complex random numbers sampled from a zero-mean Gaussian distribution with variance $1/2$ for $\ell=1,...,L$.
Quantum cumulant expansion
The many-body dynamics generated by Eq. (1) can in principle be explored by transforming the master equation for the density operator $\hat{\rho}$ into a set of coupled differential equations for the bosonic fields $\langle\hat{a}_{\ell}\rangle$. By exploiting the cyclic property of the trace, the Lindblad master equation (1) can be exactly mapped to the following set of equations,
${\rm i}\frac{\partial}{\partial t}\langle\hat{a}_{1}\rangle=-\left(\Delta+{\rm i}\gamma/2\right)\langle\hat{a}_{1}\rangle+U\langle\hat{a}_{1}^{\dagger}\hat{a}_{1}\hat{a}_{1}\rangle\ -J\langle\hat{a}_{2}\rangle+F\,,$
${\rm i}\frac{\partial}{\partial t}\langle\hat{a}_{\ell}\rangle=-\Delta\langle\hat{a}_{\ell}\rangle+U\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\hat{a}_{\ell}\rangle-J\left(\langle\hat{a}_{\ell+1}\rangle+\langle\hat{a}_{\ell-1}\rangle\right)\,,$ (33)
${\rm i}\frac{\partial}{\partial t}\langle\hat{a}_{L}\rangle=-\left(\Delta+{\rm i}\gamma/2\right)\langle\hat{a}_{L}\rangle+U\langle\hat{a}_{L}^{\dagger}\hat{a}_{L}\hat{a}_{L}\rangle-J\langle\hat{a}_{L-1}\rangle\,.$
Here, the issue is that Eqs. (Quantum cumulant expansion) are not closed, since one needs differential equations for the time evolution of $\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\hat{a}_{\ell}\rangle$. This generates an infinite BBGKY hierarchy of equations involving higher and higher-order correlations. A possible approach to obtain a closed (and solvable) set of differential equations is to truncate this infinite set by keeping only expectation values with products involving a maximum of $d$ operators. A systematic way to perform this truncation is the quantum cumulant expansion (QCE), that decomposes the expectation values of products of $D>d$ operators into a sum of expectation values of products of 1, 2, …, $d$ operators. The integer $d$ is the quantum cumulant order. A complete discussion on the truncation rules, with several examples, can be found in Ref. [77].
The QCE approach allows exploring quantum correlations beyond pure mean-field ansätze, but the set of equations for large system sizes becomes numerically intractable as we increase $d$. In particular, in the case of Eqs. (Quantum cumulant expansion), the number of cumulant equations scales as $L^{d}$. In this work, we focus on $d=2$, that captures Gaussian correlations. Calculations have been performed with the help of symbolic calculators for QCE equations available in Julia [78].
Benchmarking the TWA
We first benchmark the TWA equations (II.1) against the Lindblad master equation (1). For the parameters considered in the main manuscript, the photon number is around $60$ in the regular coherent regime. This implies that a direct solution of the Lindblad equation is difficult even in the smallest chain of $L=2$ resonators, because of the large required cutoff in the Hilbert space. We therefore exploit the Monte-Carlo wave function method described in the previous section for the solution of Eq. (1).
Expectation values of operators are obtained upon averaging over many independent quantum trajectories. In particular, photon number and its fluctuations can be respectively obtained as
[…]
In Fig. 7, we present the comparison between TWA dynamics and exact dynamics (ED) for a driven-dissipative Bose-Hubbard dimer, i.e. $L=2$. Due to the large size of the Hilbert space, we limit the exact time evolution to $t=100$ and the number of Monte Carlo trajectories to $N_{\rm traj}=500$. Furthermore, we average over the time window $[50,100]$. The same protocol is applied to the Wigner trajectories. In Fig. 7 (a), we plot the photon number $n$ for the modes $\hat{a}_{1}$ and $\hat{a}_{2}$ as a function of the drive amplitude $F$. The initial conditions are chosen as |$\ket{\psi(0)}=\ket{0}\otimes\ket{0}$ for the ED, and $\alpha_{1}(0)=\alpha_{2}(0)=0$ for the TWA dynamics. In Fig. 7 (b), we again plot $n$ for both the modes, but the initial conditions are now |$\ket{\psi(0)}=\ket{\alpha_{1}}\otimes\ket{\alpha_{2}}$ for the ED and $\alpha_{1}(0)=\alpha_{1}$, $\alpha_{2}(0)=\alpha_{2}$ for the TWA dynamics, where $\alpha_{1}$ and $\alpha_{2}$ are the steady-state values of the complex fields in each resonator computed with the TWA. In both the panels, TWA and ED show an excellent agreement. In particular, panel (b) shows that TWA and ED coincides in the NESS. In Figs. 7 (c) and (d), we repeat the same analysis of Figs. 7 (a) and (b) but for $\delta n$ instead of $n$. Again, TWA and ED show an excellent agreement.
Figure 7: Benchmark of TWA against exact dynamics (ED) for a driven-dissipative Bose-Hubbard chain described in Eq. (2) with length $L=2$. (a) Comparison between steady-state photon number $n$ obtained with TWA (colored lines) and Monte Carlo quantum trajectories (black markers) obtained by time evolving Eqs. (II.1) and the stochastic Schrödinger equation up to $t=100$. Averages have been performed over the time interval $[50,100]$ and over $500$ trajectories both for TWA and ED. The initial condition for TWA was $\alpha_{1}(0)=\alpha_{2}(0)=0$, while the initial state for ED was |$\ket{\psi(0)}=\ket{0}\otimes\ket{0}$. (b) Same analysis as in (a) but considering as initial states the coherent states $\alpha_{1}(0)=\alpha_{1}$, $\alpha_{2}(0)=\alpha_{2}$ and |$\ket{\psi(0)}=\ket{\alpha_{1}}\otimes\ket{\alpha_{1}}$, where $\alpha_{1}=\langle\hat{a}_{1}\rangle$ and $\alpha_{2}=\langle\hat{a}_{2}\rangle$ are computed with the TWA in the steady state. (c) Same analysis as in (a) but for the photon number statistics $4\delta n_{\ell}$. (d) Same analysis as in (b) but for the photon number statistics $\delta n_{\ell}$. ED and TWA show an excellent agreement in the vacuum, chaotic and high-photon number phases. The other parameters are set as in Fig. 3.
Figure 8: Comparison between TWA and the quantum cumulant expansion. (a) Photon number and (b) photon number statistics in the last resonator computed with the quantum cumulant expansion at second order as a function of the drive strength $F$ and the chain’s length $L$. Each point has been obtained upon averaging over a time window equal to $\Delta\tau=10^{3}$ in the steady state. The black dashed line indicates the boundaries of the chaotic region identified by the TWA. The other parameters are set as in Fig. 3.
We now compare TWA and the quantum cumulant expansion. In particular, we compare $n$ and $\delta n$ for the $L$-th resonator of the driven-dissipative Bose-Hubbard chain. We work with a second-order QCE approximation for which the expression of the photon number simply reads $n_{\ell}=\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle$, while $\delta n$ is now expressed as
$\delta n_{\ell}=\langle\hat{a}_{\ell}^{\dagger 2}\rangle\langle\hat{a}_{\ell}^{2}\rangle+\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle^{2}-2|\langle\hat{a}_{\ell}\rangle|^{4}\,.$ (35)
In Fig. 8 we present the comparison between the TWA and the QCE. Figs. 8 (a) and (b) show $n_{L}^{\rm ss}$ and $\delta n_{L}^{\rm ss}$ computed solving cumulant equations truncated at order two. To compute the steady-state photon number and its fluctuations at each point in the phase diagram, we let the QCE dynamics evolve to the steady state, and we subsequently averaged the dynamics over a time window equal to $\Delta\tau=10^{3}$. The black line indicates the boundaries of the chaotic region computed from TWA equations. Data show a good qualitative agreement between TWA and QCE. However, as it is already evident from a rapid comparison between Figs. 3 (b) and (c) and Fig. 8, the two methods do not show a quantitative agreement.
Additional phase diagrams
Figure 9: Saturation value of the steady-state phase OTOC, $D_{1,\ell}^{\rm ss}(\tau\to\infty)$ computed at different representative sites $\ell$ across the chain, as a function of the drive amplitude $F$ and the chain’s length $L$: (a) $\ell=1$, (b) $\ell=L/2$ , and (c) $\ell=L$ [same data as in Fig. 3 (a)]. The other parameters are set as in Fig. 3.
In this section, we provide additional plots to support the claim that the presence of quantum chaos is spread over the whole chain, and not only at the right boundary $\ell=L$ which is used in the main manuscript as a representative site. Notably, we find clear signatures of chaos in the prethermal domain, which justifies the term prethermal chaos that we used. In Fig. 9, we plot the saturation value of the phase OTOC $D_{1,\ell}^{\rm ss}(\tau\to\infty)$ as a function of the drive amplitude $F$ and the chain’s length $L$, when evaluated at three representative sites across the chain: first $\ell=1$, middle $\ell=L/2$, and last site $\ell=L$ (that was already presented in the main text).
When $\ell=1$, the saturation value never reaches 1, signaling that the first site does not reach maximal decorrelation. This is in agreement with the OTOC’s dynamics studied in Fig. 2 (a) for a specific value of $F$ and $L$. The phase diagram for $\ell=L/2$ is instead essentially identical to the one realized for $\ell=L$. This analysis shows that the prethermal and thermal domains are equally chaotic.
Comparison with the uniformly driven-dissipative 1D chain
Figure 10: Comparison between the lattices with boundary and uniform drive and dissipation for $L=10$. (a) Photon number in the last site as a function of $F$, (b) photon number statistics in the last site as a function of $F$, (c) saturation value of the steady-state OTOC in the last site as a function of $F$. The results are obtained upon averaging over $10^{3}$ independent Wigner trajectories. The other parameters are set as in Fig. 3.
Here, we compare the boundary-driven, boundary-dissipative Bose-Hubbard lattice (1), for which we have found a wide chaotic region in parameter space, with the uniformly-driven uniformly-dissipative Bose-Hubbard chain. The Hamiltonian of the latter model is given by
$\hat{H}=\sum_{\ell=1}^{L}\left[-\Delta\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}+\frac{1}{2}U\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\hat{a}_{\ell}+F(\hat{a}_{\ell}^{\dagger}+\hat{a}_{\ell})\right]-J\sum_{\ell=1}^{L-1}\left(\hat{a}_{\ell+1}^{\dagger}\hat{a}_{\ell}+\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell+1}\right),$ (36)
while the Lindblad master equation reads
$\frac{\partial\hat{\rho}}{\partial t}=-{\rm i}[\hat{H},\hat{\rho}]+\sum_{\ell=1}^{L}\gamma\left(\hat{a}_{\ell}\hat{\rho}\hat{a}_{\ell}^{\dagger}-\frac{1}{2}\anticommutator{\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}}{\hat{\rho}}\right).$ (37)
In Fig. 10, we compare the boundary-driven boundary-dissipative chain studied in this paper and the uniformly-driven uniformly-dissipative chain, studied, e.g., in Ref. [40], showing that quantum chaos is a feature peculiar of boundary drive and dissipation mechanisms. We focus on the last cavities of the two chains in the steady-state regime. In Fig. 10 (a), we show that the photon number $n_{L}^{\rm ss}$ exhibits a similar behavior with an abrupt jump around the same critical drive amplitude $F$. In Fig. 10 (b), we present the photon number statistics $\delta n_{L}^{\rm ss}$. In the case of the uniformly-driven uniformly-dissipative chain, large fluctuations are concentrated at the value of $F$ where the jump occurs, as expected. In contrast, in the case of boundary drive and dissipation, large fluctuations are spread over a much larger region. In Fig. 10 (c), we plot the saturation value of the steady-state semiclassical phase OTOC defined in Eq. (5), showing final evidence that only the boundary-driven, boundary-dissipative lattice hosts quantum chaos, while the dynamics in the uniformly-driven dissipative chain is regular.
Modeling the local state of the chaotic Bose-Hubbard chain with a U(1)-symmetric impurity model
Models and effective temperature
We consider two single-site driven-dissipative impurity models that describe U(1)-symmetric quantum states exhibiting, across their parameter space, thermal features and population inversion. Population inversion is achieved when higher energy states are more populated than lower energy states, in contrast to a thermal population, and requires an external energy source [79, 67, 80]. Population inversion is an essential ingredient to obtain lasing states, where stimulated emission dominates over absorption [81, 82]. The Hamiltonian of both the models is taken as the local Hamiltonian of the Bose-Hubbard chain in Eq. (2), $\hat{H}_{\rm imp}=\omega_{0}\hat{a}^{\dagger}\hat{a}+U\hat{a}^{\dagger 2}\hat{a}^{2}/2$. The first impurity model, which we refer as the 2-photon decay model, is defined by the set of Lindblad dissipators described in the main text,
$\hat{L}_{\uparrow}=\sqrt{\gamma^{\uparrow}}\hat{a}^{\dagger},\qquad\hat{L}_{\downarrow}=\sqrt{\gamma^{\downarrow}}\hat{a},\qquad\hat{L}_{\phi}=\sqrt{\gamma^{\phi}}\hat{a}^{\dagger}\hat{a},\qquad\hat{L}_{\rm s}=\sqrt{\gamma^{\rm s}}\hat{a}^{2}\,.$ (38)
The non-negatives parameters $\gamma^{\uparrow}$, $\gamma^{\downarrow}$, $\gamma^{\phi}$, and $\gamma^{\rm s}$ are effective rates of incoherent pumping, decay, dephasing, and 2-photon decay. The second impurity model is the generalized Scully-Lamb model, introduced in Ref. [65] and characterized by the Lindblad dissipators
$\hat{L}_{1}=\hat{a}^{\dagger}(\gamma^{\uparrow}-\mathcal{S}\hat{a}\hat{a}^{\dagger})/\sqrt{\gamma^{\uparrow}},\qquad\hat{L}_{\phi}=\sqrt{\gamma^{\phi}}\hat{a}\hat{a}^{\dagger},\qquad\hat{L}_{\downarrow}=\sqrt{\gamma^{\downarrow}}\hat{a}.$ (39)
The non-negatives parameters $\gamma^{\uparrow}$, $\gamma^{\downarrow}$, $\gamma^{\phi}$, and $\mathcal{S}$ are effective rates of incoherent pumping, decay, dephasing, and saturation. If $\gamma^{\phi}=3\mathcal{S}/2$ the above model reduces to the standard Scully-Lamb model of Ref. [66]. The Scully-Lamb model is valid if [83]
$\gamma^{\uparrow}\sim\mathcal{O}(\gamma^{\downarrow}),\qquad\mathcal{S}\langle\hat{a}\hat{a}^{\dagger}\rangle\ll\gamma^{\uparrow},$ (40)
while for the first model one requires $\gamma^{\rm s}\langle\hat{a}\hat{a}^{\dagger}\rangle\ll\gamma^{\uparrow}$. As discussed in the main text, the choice of $\hat{H}_{\rm imp}$ and $\gamma^{\phi}$ does not affect the steady-state properties, including the steady-state Wigner function $W(\alpha,\alpha^{*})$. In the steady state, they have an effect only on time-correlation functions. In particular, by imposing $\gamma^{\phi}=3\mathcal{S}/2$ one finds oscillating time-correlation functions with a usually very small decay rate (that is $23\mathcal{S}/$ 2 which is by assumption much smaller than $2\gamma^{\uparrow}/\langle\hat{a}\hat{a}^{\dagger}\rangle$). If instead $\gamma^{\phi}\gg\mathcal{S}$ one obtains the generalized version of the Scully-Lamb model exhibiting rapidly decaying time-correlation functions in the lasing phase. A similar phenomenology occurs in the 2-photon decay impurity model, where the phase coherence is suppressed if $\gamma^{\phi}\gg\gamma^{\rm s}$.
While the two impurity models above exhibits very similar steady-state properties, a finite (inverse) effective temperature $\beta^{\rm eff}$ can only be simply associated with the 2-photon decay model. Indeed, the set of Lindblad operators in Eq. (38) can be easily interpreted as coupling to a set of thermal reservoirs. $\hat{L}_{\rm s}$ corresponds to a zero-temperature 2-photon reservoir. $\hat{L}_{\uparrow}$ and $\hat{L}_{\downarrow}$ correspond to finite-temperature 1-photon reservoir at a temperature set by the detailed balance
$\frac{\gamma^{\uparrow}}{\gamma^{\downarrow}}=:{\rm e}^{-\beta^{\rm eff}(\omega_{0}-\mu)}=:{\rm e}^{\beta^{\rm eff}\mu^{\rm eff}},\quad\textit{i.e.}\quad\beta^{\rm eff}\mu^{\rm eff}:=\log\left({\gamma_{\uparrow}}\big{/}{\gamma_{\downarrow}}\right),$ (41)
where we neglected $U\ll\omega_{0}$ since, in typical circuit QED setups, the resonator frequency is of the order of GHz while the Kerr nonlinearity ranges from kHz to MHz [44]. The dephasing operator $\hat{L}_{\phi}$ is self-adjoint and corresponds to infinite temperature.
In contrast, inspecting of the set of Lindblad operators of the generalized Scully-Lamb model given in Eq. (39), one sees that $\hat{L}_{1}$ is not the adjoint of $\hat{L}_{\downarrow}$ because of the presence of the nonlinear saturation $\mathcal{S}$. Thus, the balance between $\hat{L}_{1}$ and $\hat{L}_{\downarrow}$ cannot be used to define an effective temperature.
Fitting procedure for the Wigner function
Figure 11: Fit of the $\ell$-th site Wigner function of the driven-dissipative bosonic chain described by Eqs. (2) and (1) for a chain with length $L=400$. In panel $(a)$, we show the $L^{2}$ norm in Eq. (42) between the $L$-th site Wigner function and the Scully-Lamb fitted Wigner function for one site each ten along the chain (green solid line). Additionally, we present the $L^{2}$ norms associated with the 2-photon impurity ansatz in Eq. (38) (purple dashed line) and the Gibbs state in Eq. (45) (black dotted line). Panels (b-e) shows the comparison between two cuts of $W(\alpha,\alpha^{*})$ [the one at $\textrm{Re}(\alpha)=0$ (blue curve) and the one at the one at $\textrm{Im}(\alpha)=0$ (orange curve)] and the fitted Wigner function (black dashed line). The values of $\ell$ are (b) $\ell=1$, (c) $\ell=200$, (d) $\ell=400$.
[…]
$L^{2}(W_{\ell},W^{\textrm{fit}}_{\ell})=\left[\int_{\mathcal{A}}{\rm d}\alpha{\rm d}\alpha^{*}\,|W_{\ell}(\alpha,\alpha^{*})-W^{\textrm{fit}}_{\ell}(\alpha,\alpha^{*})|^{2}\right]^{1/2}.$ (42)
In Fig. 11, we show the results of the fitting procedure corresponding to Fig. 5. In Fig. 11 (a), we plot the $L^{2}$ norm as a function of the index site $\ell$ for the chain with length $L=400$, using both the 2-photon decay and the generalized Scully-Lamb model as single-site ansätze. The results are compared with the $L^{2}$ norm corresponding to the fit of the thermal ansatz discussed in the following section. When $\ell\leq 50$ the fitting procedure does not give reliable results. In the bulk and in the right tail of the chain, instead, the local Wigner function is completely captured by both the single-site impurity models. In Figs. 11 (b-d) we show the comparison between $W_{\ell}(\alpha,\alpha^{*})$ and $W^{\rm fit}_{\ell}(\alpha,\alpha^{*})$. For $\ell=1$ [c.f. Fig. 11 (b)], the local Wigner function $W_{\ell}(\alpha,\alpha^{*})$ is not captured by the single-site ansatz. We conclude that within the left tail of the chain, in proximity of the coherent drive, different cavities remain correlated and thus a single-site description is not sufficient. For $\ell=200,400$ [c.f. Fig. 11 (c) and (d)] we see how $W_{\ell}(\alpha,\alpha^{*})$ matches with $W^{\rm fit}_{\ell}(\alpha,\alpha^{*})$.
Classical auto-correlation function
Figure 12: Classical auto-correlation function in the prethermal phase of a chaotic Bose-Hubbard chain with length $L=200$. We plot $C_{\ell}(\tau)$ computed according to Eq. (43) for (a) $\ell=30$, (b) $\ell=60$, and (c) $\ell=100$. Results have been obtained from a single long Wigner trajectory in the steady state. The black dashed line indicates the fit performed according to Eq. (44). The drive strength is fixed to $F=7.5$ and the other parameters are set as in Fig. 3.
Here, we show that, in the prethermal phase of the chaotic Bose-Hubbard chain, the dephasing rate $\gamma^{\phi}$ is much larger than the 2-photon decay rate $\gamma^{\rm s}$ or the saturation $\mathcal{S}$, hence establishing the absence of long-lived phase coherence typical of lasing states. As discussed above, $\gamma^{\phi}$ cannot be extracted from the local steady-state Wigner function $W^{\rm ss}_{\ell}(\alpha,\alpha^{*})$. Here, we extract a rough estimate of $\gamma^{\phi}$ from two-time correlation functions in the prethermal domain in an $L=200$ chaotic chain by fitting them to a simple ansatz.
We consider the classical steady-state auto-correlation function defined as
$C_{\ell}(\tau):=\textrm{Re}\lim_{t,T\to+\infty}\frac{1}{T}\int_{t}^{t+T}{\rm d}t^{\prime}\,\,\frac{\langle\hat{a}_{\ell}^{\dagger}(t^{\prime}+\tau)\rangle\langle\hat{a}_{\ell}(t^{\prime})\rangle}{|\langle\hat{a}_{\ell}(t^{\prime})\rangle|^{2}},$ (43)
that can be computed over a single long Wigner trajectory in the steady state, due to the ergodic nature of steady-state trajectories [84].
For the Scully-Lamb model, i.e. $\gamma^{\phi}=3\mathcal{S}/2$, time-correlation functions exhibit a slow oscillatory decay controlled by the rate $\mathcal{S}$. For the generalized Scully-Lamb model, the much faster decay of time-correlation functions is controlled by $\gamma^{\phi}\gg\mathcal{S}$ [65]. A similar phenomenology occurs in the 2-photon decay impurity model. In Fig. 12, we plot $C_{\ell}(\tau)$ for a chain of length $L=200$ with a drive amplitude $F=7.5$ and at three sites within the prethermal domain: (a) $\ell=30$, (b) $\ell=60$, and (c) $\ell=100$. The auto-correlation function exhibits a rapid decay towards zero, with an additional weak oscillation. To quantify the decay rate, we fit the initial decay of $C_{\ell}(\tau)$ with
$C_{\ell}^{\rm fit}(\tau)=\sum_{j=1}^{2}A_{\ell,j}{\rm e}^{-\Gamma_{\ell,j}\tau}\cos(\Omega_{\ell,j}\tau).$ (44)
Here, $A_{\ell,j}$ are real-valued amplitudes, $\Gamma_{\ell,j}\geq 0$ are decaying rates with magnitudes on the order of $\gamma^{\phi}$ [65], while the frequencies $\Omega_{\ell,j}$ capture possible oscillations in the early times. This simple ansatz successfully captures the auto-correlation function in the prethermal phase, as shown in Fig. 12 (black dashed lines). In the following table, we compare the resulting decay rates $\Gamma_{1}$ and $\Gamma_{2}$ with the saturation parameters obtained from the fit of $W^{\rm ss}_{\ell}(\alpha,\alpha^{*})$.
$\Gamma_{1}$ $\Gamma_{2}$ $\gamma^{\rm s}$ $\mathcal{S}$
$\ell=30$     0.676     0.358     0.0012     0.0011
$\ell=60$     0.670     0.678     0.0010     0.0009
$\ell=100$     0.954     0.954     0.0007     0.0007
In the three cases, the exponential decay rate of $C_{\ell}(\tau)$ is much larger than $\gamma^{\rm s}$ and $\mathcal{S}$, indicating the presence of a large dephasing rate $\gamma^{\phi}$, that is included in the 2-photon decay and generalized Scully-Lamb models. This analysis shows that the prethermal domain is not a laser since the latter would exhibit a significant phase coherence in time. Note that this distinction with a laser relies on the analysis of dynamical properties of the prethermal domain that go beyond the characterization of its static properties. This motivates our choice to describe the local physics of the chain in terms of a driven-dissipative impurity model rather than a Gibbs state ansatz (see the following section).
Modeling the local state of the chaotic Bose-Hubbard chain with a Gibbs state
Figure 13: Comparison between two cuts of $W(\alpha,\alpha^{*})$ [the one at $\textrm{Re}(\alpha)=0$ (blue curve) and the one at the one at $\textrm{Im}(\alpha)=0$ (orange curve)] and the fitted Wigner functions with the Gibbs state in Eq. (45), by imposing $U=0$ (green dashed line) and $U=0.1$ (black dash dotted line). Data are plotted in log scale on the y axis for a better visualization of the differences between the two fits. The drive strength is fixed to $F=7.5$ and the other parameters are set as in Fig. 3.
As explained in the main text, the local statics can also be captured by a Gibbs state reading
$\hat{\rho}^{\rm eq}_{\ell}=\frac{\exp[-\beta_{\ell}^{\rm eff}(\hat{H}_{\rm imp}-\mu_{\ell}\,\hat{a}^{\dagger}a)]}{\operatorname{Tr}\left(\exp[-\beta_{\ell}^{\rm eff}(\hat{H}_{\rm imp}-\mu_{\ell}\,\hat{a}^{\dagger}a)]\right)}\,,$ (45)
where the impurity Hamiltonian $\hat{H}_{\rm imp}$ is given by Eq. (8), $\beta^{\rm eff}_{\ell}$ is the effective temperature and $\mu_{\ell}$ is the chemical potential at the $\ell$-th site. By reabsorbing all the quadratic contributions into the effective chemical potential $\mu_{\ell}^{\rm eff}:=\mu_{\ell}-(\omega_{0}-U/2)$, we find that the local Wigner functions can simply be described by the two parameters $\beta_{\ell}^{\rm eff}$ and $\mu_{\ell}^{\rm eff}$. Note that, contrary to the driven-dissipative impurity models, the precise content of $\hat{H}_{\rm imp}$ is relevant for the fitting procedure to the Gibbs state. In particular, we fix $U=0.1$ [the value we set in all the numerical simulations of Eq. (2) presented in this work]. The $L^{2}$ norm associated to this fitting procedure is plotted in Fig. 11 (a). The result shows that the Gibbs state in Eq. (45) is successful in describing the local physics of the chain in the prethermal and thermal domains.
In Fig. 5 (a), we presented the product $\beta_{\ell}^{\rm eff}\mu_{\ell}^{\rm eff}$ for the three models considered in the paper, namely the 2-photon decay model in Eq. (38), the Scully-Lamb model in Eq. (39), and the Gibbs state in Eq. (45). All the proposed single-site ansätze give qualitatively similar results, thus validating the robustness of our local approach in capturing the features of $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$. A more sophisticated study of space and time correlation functions, and their characterization with multiple-site impurities, should be addressed in future studies.
Finally, we show that taking into account the Kerr nonlinearity $U$ is important to the characterization of the local state in the thermal phase. In Fig. 13, we plot the Wigner function for $\ell=400$ at $\textrm{Im}(\alpha)=0$ and $\textrm{Re}(\alpha)=0$. We first perform the fit to the Gibbs state by imposing $U=0$ (black dashed line), which coincides with a Gaussian fit with a single parameter, $\beta_{\ell}^{\rm eff}\mu_{\ell}^{\rm eff}$. In this case, the fit does not reproduce the local Wigner functions. Instead, when we impose $U=0.1$ (gray dash-dotted line), the two-parameters ansatz perfectly reproduces the shape of $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$.